If you take sodium iodide (10.7 g) and dissolve it to 250 mL, what is the concentration?
Answer #1

molarity = weight of substance*1000/gram molar mass*volume in ml

weight of substance = 10.7gm

gram molar mass of NaI = 23 + 127 = 150gm/mole

volume of solution = 250ml

molarity = 10.7*1000/150*250 = 0.285M

concentration = 0.285M